OpenStudy (anonymous):

what is the vertex form for:y=-2x^2+12x-18 ?

jimthompson5910 (jim_thompson5910):

x = -b/(2a) x = -12/(2(-2)) x = 3 Plug this into the equation to find y. Tell me what you get.

OpenStudy (anonymous):

I suggest to reorder the equation

OpenStudy (anonymous):

by completing squares